# Aioli: A unified optimization framework for language model data mixing

### v1

@@ Abstract
Language model performance depends on identifying the optimal mixture of data groups to train on (e.g., law, code, math). Prior work has proposed a diverse set of methods to efficiently learn mixture proportions, ranging from fitting regression models over training runs to dynamically updating proportions throughout training. Surprisingly, we find that no existing method consistently outperforms a simple stratified sampling baseline in terms of average test perplexity per group. In this paper, we study the cause of this inconsistency by unifying existing methods into a standard optimization framework. We show that all methods set proportions to minimize total loss, subject to a method-specific mixing law—an assumption on how loss is a function of mixture proportions. We find that existing parameterizations of mixing laws can express the true loss-proportion relationship empirically, but the methods themselves often set the mixing law parameters inaccurately, resulting in poor and inconsistent performance. Finally, we leverage the insights from our framework to derive a new online method named Aioli, which directly estimates the mixing law parameters throughout training and uses them to dynamically adjust proportions. Empirically, Aioli outperforms stratified sampling on 6 out of 6 datasets by an average of 0.28 test perplexity points, whereas existing methods fail to consistently beat stratified sampling, doing up to 6.9 points worse. Moreover, in a practical setting where proportions are learned on shorter runs due to computational constraints, Aioli can dynamically adjust these proportions over the full training run, consistently improving performance over existing methods by up to 12.01 test perplexity points.
@@ 1 Introduction
It is critical to determine what data to train on for a language model (LM) to acquire a range of capabilities, from generating code to understanding scientific literature and conversing with users [3, 39, 34]. To achieve this, practitioners mix data from various groups (such as code files, scientific papers, and chat logs) in specific proportions to compose an overall training dataset—a procedure known as data mixing. Identifying the optimal mixture proportions is critical to LLM performance. However, a brute-force trial-and-error search over the proportions is computationally expensive, requiring many training runs.
Recent work introduces two types of data mixing algorithms that learn mixture proportions: offline and online methods. Offline methods conduct multiple training runs with varying proportions, fit a regression model to predict performance, and use this model to determine the optimal static mixture [72, 37]. Online methods adjust the mixture proportions dynamically throughout training using information from the model, such as its loss and gradients [13, 21, 70, 2]. These mixing methods require at least one training run to learn the proportions but are more efficient than a brute-force search.
Given the wide range of methods available, it is important to determine which ones are effective. However, when we evaluated existing methods, we found that no method consistently outperformed stratified sampling—a simple baseline that uniformly mixes groups and requires zero extra training runs—across all sets of data groups in terms of average test perplexity (Table 2). This surprising outcome suggests that all existing methods suffer from some common weaknesses. To make progress in data mixing, we identify three objectives for our work: 1) improve our understanding of the underlying assumptions of existing methods, 2) assess the fidelity of these assumptions in practice to better understand performance, and 3) apply our insights to develop principled new data mixing methods.
In this paper, we improve our understanding of data mixing methods by showing that many existing methods can be expressed in a unified optimization framework, which we call Linear Mixing Optimization (LMO) (Section 3). These methods solve an optimization problem that sets proportions to minimize the average loss per data group, subject to a method-specific mixing law—a particular assumption relating loss per group and mixture weights. We find that all current mixing laws share the same parameterization: for training round $t$ from $1$ to $T$,
@@
Applying the LMO framework, we test the fidelity of existing methods assumptions, examining if they hold in practice (Section 4). Both the log-linear static and linear dynamic parameterizations capture the true loss-proportion relationship across pretraining datasets, achieving an average of 0.0005 MSE and 0.969 $R^{2}$. We then show that although existing mixing laws are well-specified, methods can set their parameters ($A^{t}$) inaccurately, causing poor performance. We compare each method’s parameters to the optimal parameters, which we approximate by fitting the mixing laws to training runs. We find that the method’s parameters can differ significantly from the optimal parameters, and the extent of these deviations is correlated with method performance relative to stratified sampling (Figure 3), helping explain our initial observations. Finally, we validate the assumptions made in solving the optimization problem, finding that the greedy approximation in online methods is a reasonable proxy for the full objective. Our analysis shows that existing methods’ parameterizations and solving strategies are of high fidelity, but their parameters are not.
To validate these insights, we develop Aioli, a simple new online data mixing method derived from the LMO framework (Section 5). Similar to existing online methods, Aioli uses exponentiated gradient descent to minimize the average loss per group at each timestep subject to a linear dynamic mixing law, which we have empirically shown to be well-specified. Unlike existing online methods, Aioli directly estimates the parameters $A^{t}$ from the current training run by fitting the mixing law on the history of losses and dynamic mixture proportions so far. Aioli is thus able to dynamically adjust proportions without requiring any extra training runs.
We evaluate Aioli in two settings by training 160M parameter models on various combinations of data sources from SlimPajama [53] (Section 6). First, we compare Aioli to existing data mixing methods and find that Aioli consistently outperforms stratified sampling on all 6 datasets, by an average of $0.280$ and up to $0.439$ points in test perplexity. On the other hand, existing data mixing methods do worse than stratified on at least one dataset by up to 6.9 perplexity points, despite using extra training runs. As we expect, the parameters of Aioli are also more consistently close to the optimal parameters (Figure 2). Second, we consider a scenario with limited additional computational resources, in which practitioners cannot run experiments for learning mixture proportions for the full training duration. In this setting, mixture proportions learned on a shorter run may not perform well on the longer final run. We find that using Aioli to dynamically adjust these learned proportions throughout the final training run can improve performance by an average of 1.202 perplexity points in 28 out of 30 cases, compared to using the learned proportions directly.
@@ 2 Problem Setup
Data & Mixing.
During training, we show the model a total of $N$ examples from $D_{\text{train}}$ over $S$ training steps. To express how data proportions can change throughout training, we divide training into $T$ equal rounds. Each round $t$ uses a mixture proportion from the probability simplex: $p^{t}=\left[p_{1}^{t},\ldots,p_{m}^{t}\right]\in\triangle^{m}$. Static mixtures use only a single round ($T=1$): $\bm{p}=\left(p^{1}\right)$, while dynamic mixtures use several ($T>1$): $\bm{p}=\left(p^{1},\ldots,p^{T}\right)$.
Model & Loss.
Let $f(\bm{p},t)$ refer to the language model (LM), $f$, at the beginning of round $t$ where the model has been trained on data sampled using mixture proportions $p^{1},\cdots,p^{t-1}$ so far. Given a model $f$, we can compute its loss on each group using the training data, $L_{\text{train}}(f)=\left(L_{\text{train},1}(f),\dots,L_{\text{train},m}(f)\right)$, and similarly using the validation data, $L_{\text{val}}(f)$, and test data, $L_{\text{test}}(f)$. In this notation, the loss at the end of training can be expressed as $L_{(\cdot)}(f(\bm{p},T+1))$. When the $f$ being referred to is obvious, we simply write $L_{(\cdot)}^{t}(\bm{p})$, and for static mixtures we drop the superscript: $L_{(\cdot)}(\bm{p})$.
Data Mixing Problem.
Given a set of data groups, an LM $f$ to train for $S$ steps with $N$ samples, and $T$ rounds of training (i.e., determining whether we use static or dynamic proportions), we aim to determine the $\bm{p}$ that minimizes the total test loss across groups: $\underset{\bm{p}\in\triangle^{T\times m}}{\text{minimize}}\sum_{i=1}^{m}L_{\text{test},i}^{T+1}(\bm{p}).$
This objective aims to produce a trained model that does well on many data groups, which can serve as a proxy for downstream performance. However, without assuming additional structure on $L_{\text{test}}^{T+1}(\bm{p})$, this problem can only be solved with a brute-force search over $\bm{p}$, which requires training many different models. Existing methods overcome this by imposing an implicit constraint on the problem, thereby setting $\bm{p}$ without searching. In the next section, our LMO framework unifies many existing methods using a single explicit constraint.
@@ 3 A Unified Optimization Framework for Data Mixing
We introduce the LMO framework by stating the general optimization problem (Section 3.1). Then, we explain how this framework can express several existing methods (Section 3.23.3), with a summary of our insights regarding these methods in Section 3.3.3.
3.1 Linear Mixing Optimization (LMO)
The LMO framework consists of an optimization problem that is equivalent to the data mixing problem (Section 2), subject to an additional constraint:
@@ 3.1 Linear Mixing Optimization (LMO)
There are three components of this optimization problem that need to be specified to yield a way to set $\bm{p}$: a) the parameterization of the mixing law, defined by $T$ and $\sigma$; b) the values of the parameters $A^{t},b^{t},$ and $c^{t}$; and c) how to solve the problem. We express existing methods in LMO by specifying these components.
3.2 Preliminaries for expressing methods in the LMO framework
We discuss preliminaries before presenting existing methods and explaining how they can be expressed in the LMO framework. First, we formally define what it means for a method to be expressed in the LMO framework. Then, we present a result that allows us to convert between linear dynamic mixing laws ($T>1$, $\sigma=\text{Id}$) and a way to set $\bm{p}$, which we will to use to express online methods in our framework in Section 3.3.
@@ 3.2 Preliminaries for expressing methods in the LMO framework
A data mixing method can be expressed in the LMO framework if the way it sets proportions $\bm{p}$ and trains model $f$ in terms of $\bm{p}$ can be equivalently constructed by specifying a mixing law parameterization, parameters, and way of solving the LMO optimization problem.
This definition allows us to cast existing methods as a way of solving the LMO optimization problem based on how they set $\bm{p}$ and train according to $\bm{p}$, even if the methods themselves are not originally designed to minimize average validation loss.
Converting mixing laws into update rules.
When $T>1$, a natural way to solve the LMO optimization problem is via exponentiated gradient descent (EGD) [31, 5], which updates $p^{t}$ greedily while ensuring that it remains on the probability simplex. The following lemma presents the EGD update rule for the LMO optimization problem when $\sigma=\text{Id}$.
@@
This lemma shows how to adjust $p^{t}$ dynamically to solve the LMO optimization problem. Notably, this update rule is defined in terms of the mixing law parameters, $A^{t}$ and $b^{t}$. This gives us a way to convert between how a method sets $\bm{p}$ and the implicit assumption it makes in its mixing law.
@@
Table 1: Summary of how existing methods and Aioli are expressed in the LMO framework.
3.3 Existing methods
We discuss four existing data mixing methods and express them as specific instances of the LMO framework. A summmary of our insights is provided in Section 3.3.3 and Table 1. In Appendix B.1, we comment on how several other online and offline data mixing methods are related to our framework, and all proofs for this section are in Appendix B.2.
@@ 3.3.1 Offline methods
Data Mixing Laws (DML).
Ye et al. [72] propose an offline method using a static mixing law ($T=1$): $L_{\text{val},i}(\bm{p})=c_{i}+b_{i}\exp(\sum_{j=1}^{m}-A_{ij}p_{j})$ for $i\in[m]$, with $A,b,c$ learned by sweeping training runs over static proportions (at least $m+1$ runs to avoid being underdetermined). Their method selects the proportion that minimizes the predicted validation loss. This law can be derived from (2) with $\sigma=\exp$, showing that LMO with a) log-linear static mixing law, b) fitted parameters, and c) direct computation of $\bm{p}$ can express DML.
@@ 3.3.2 Online Methods
We provide a colloquial description and an algorithmic description of three online methods. Then, in Theorem 1, we demonstrate how they all are expressed in LMO using a linear dynamic mixing law, the EGD update rule, and method-specific mixing law parameters.
Skill-It.
Chen et al. [13] is an online method motivated by curriculum learning that dynamically adjusts mixture proportions. Data group interactions are expressed in a “skills graph,” where each edge denotes how much the loss on one group changes when trained on another. The skills graph is learned in advance using $m$ additional training runs and is then used to update proportions $p^{t}$ throughout training.
@@
DoReMi.
Xie et al. [70] is an online method that applies ideas from distributionally robust optimization to data mixing, where the training objective minimizes the worst-group excess loss over a model trained with stratified sampling. $p^{t}$ is updated dynamically to minimize this excess loss and then averaged for the final run. DoReMi requires two additional runs to learn a static $\bm{p}$.
@@
DoGE.
Fan et al. [21] is an online method that solves a bi-level optimization problem in which $p^{t}$ is updated to minimize the average training loss at each step. By using a first-order Taylor approximation of the training loss, $p^{t}$ is updated using the gradient dot products across data groups. The dynamic proportions are then averaged for the final run. DoGE requires one additional run to learn a static $\bm{p}$.
@@
Framework expression.
All three online methods use an update rule $p_{j}^{t+1}\propto p_{j}^{t}\exp(\cdot)$, which is similar to (3). This provides intuition for our main theorem, which expresses these methods in LMO.
@@ 3.3.3 Summary of LMO Framework Insights
Table 1 summarizes how existing methods are expressed in the LMO framework. LMO reveals the assumptions each method makes through how the components of the framework are specified. First, all mixing laws are either linear or log-linear. Second, the mixing laws differ in the values of the parameters used. For example, Skill-It’s $A^{t}$ is the current loss times a matrix learned from training on static proportions, while DoReMi’s $A^{t}$ is diagonal. Third, offline mixing methods solve for $\bm{p}$ directly while online mixing methods use EGD, which uses a greedy approximation. In the next section, we study if each of these assumptions holds well in practice.

### v2

@@ Abstract
Language model performance depends on identifying the optimal mixture of data groups to train on (e.g., law, code, math). Prior work has proposed a diverse set of methods to efficiently learn mixture proportions, ranging from fitting regression models over training runs to dynamically updating proportions throughout training. Surprisingly, we find that no existing method consistently outperforms a simple stratified sampling baseline in terms of average test perplexity. To understand this inconsistency, we unify existing methods into a standard framework, showing they are equivalent to solving a common optimization problem: minimize average loss subject to a method-specific mixing law—an implicit assumption on the relationship between loss and mixture proportions. This framework suggests that measuring the fidelity of a method’s mixing law can offer insights into its performance. Empirically, we find that existing methods set their mixing law parameters inaccurately, resulting in the inconsistent mixing performance we observe. Using this insight, we derive a new online method named Aioli, which directly estimates the mixing law parameters throughout training and uses them to dynamically adjust proportions. Aioli outperforms stratified sampling on 6 out of 6 datasets by an average of 0.27 test perplexity points, whereas existing methods fail to consistently beat stratified sampling, doing up to 6.9 points worse. Moreover, in a practical setting where proportions are learned on shorter runs due to computational constraints, Aioli can dynamically adjust these proportions over the full training run, consistently improving performance over existing methods by up to 12.012 test perplexity points.
**footnotetext: Equal contribution. Contact: mfchen@stanford.edu, michael.hu@nyu.edu
@@ 1 Introduction
It is important to determine what data to train on for a language model (LM) to acquire a range of capabilities, from generating code to understanding scientific literature and conversing with users [3, 39, 34]. To achieve this, practitioners mix data from various groups (such as code files, scientific papers, and chat logs) in specific proportions to compose an overall training dataset—a procedure known as data mixing. Identifying the optimal mixture proportions is critical to LLM performance. However, a brute-force trial-and-error search over the proportions is computationally expensive, requiring many training runs.
Recent work introduces two types of data mixing algorithms that learn mixture proportions: offline and online methods. Offline methods conduct multiple training runs with varying proportions, fit a regression model to predict performance, and use this model to determine the optimal static mixture [72, 37]. Online methods adjust the mixture proportions dynamically throughout training using information from the model, such as its loss and gradients [14, 22, 70, 2]. All mixing methods require at least one training run to learn the proportions but are more efficient than a brute-force search.
Given the wide range of methods available, it is important to determine which ones are effective. However, when we evaluated existing methods, we found that no method consistently outperformed stratified sampling—a simple baseline that uniformly mixes groups and requires zero extra training runs—across all sets of data groups in terms of average test perplexity (Table 2). This surprising outcome suggests that all existing methods suffer from some common weaknesses. To make progress in data mixing, we identify three objectives: 1) improve our understanding of the underlying assumptions of existing methods, 2) assess the fidelity of these assumptions in practice to better understand performance, and 3) apply our insights to develop principled new data mixing methods.
In this paper, we improve our understanding of data mixing methods by showing that many existing methods can be expressed in a unified optimization framework, which we call Linear Mixing Optimization (LMO) (Section 3). These methods are equivalent to solving an optimization problem that sets proportions to minimize the average loss per data group, subject to an implicit method-dependent mixing law—an assumption relating loss per group and mixture proportions. We find that all current mixing laws share the same parameterization: for training round $t$ from $1$ to $T$,
@@
Applying the LMO framework, we test the fidelity of existing methods’ assumptions, examining if they hold in practice (Section 4). Both the log-linear static and linear dynamic parameterizations capture the true loss-proportion relationship across datasets, achieving an average of 0.0005 MSE and 0.969 $R^{2}$. We then show that although existing mixing laws are well-specified, methods can set their parameters ($A^{t}$) inaccurately, causing poor performance. We compare each method’s parameters to the optimal parameters, which we approximate by fitting the mixing laws to training runs. We find that the method’s parameters can differ significantly from the optimal parameters, and the extent of these deviations is correlated with method performance relative to stratified sampling (Figure 3), helping explain our initial observations. Finally, we validate the assumptions used in solving the optimization problem, finding that the greedy approximation in online methods is a reasonable proxy for the full objective. Our analysis shows that existing methods’ parameterizations and solving strategies are of high fidelity, but their parameters are not.
To validate these insights, we develop Aioli, a simple new online data mixing method derived from the LMO framework (Section 5). Unlike existing online methods, Aioli directly estimates the parameters $A^{t}$ from the current training run by fitting the mixing law on the history of losses and dynamic mixture proportions so far. Aioli is thus able to dynamically adjust proportions without requiring any extra training runs.
We evaluate Aioli in two settings by training 160M models on various combinations of data sources from SlimPajama [54] (Section 6). First, we compare Aioli to existing data mixing methods and find that Aioli consistently outperforms stratified sampling on all 6 datasets, by an average of $0.274$ and up to $0.439$ points in test perplexity. On the other hand, existing data mixing methods do worse than stratified on at least one dataset by up to 6.9 perplexity points, despite using extra training runs. As we expect, the parameters of Aioli are also more consistently close to the optimal parameters (Figure 2). Second, we consider a scenario with limited additional computational resources, in which practitioners cannot run experiments for learning mixture proportions for the full training duration. In this setting, mixture proportions learned on a shorter run may not perform well on the longer final run. We find that using Aioli to dynamically adjust these learned proportions throughout the final training run can improve performance by an average of 1.202 perplexity points in 28 out of 30 cases, compared to using the learned proportions directly.
@@ 2 Problem Setup
Data & Mixing. During training, we show the model a total of $N$ examples from $D_{\text{train}}$ over $S$ training steps. To express how data proportions can change throughout training, we divide training into $T$ equal rounds. Each round $t$ uses a mixture proportion from the probability simplex: $p^{t}=\left[p_{1}^{t},\ldots,p_{m}^{t}\right]\in\triangle^{m}$. Static mixtures use only a single round ($T=1$): $\bm{p}=\left(p^{1}\right)$, while dynamic mixtures use several ($T>1$): $\bm{p}=\left(p^{1},\ldots,p^{T}\right)$.
Model & Loss. Let $f(\bm{p},t)$ refer to the language model, $f$, at the beginning of round $t$ where the model has been trained on data sampled using mixture proportions $p^{1},\cdots,p^{t-1}$ so far. Given a model $f$, we can compute its loss on each group using the training data, $L_{\text{train}}(f)=\left(L_{\text{train},1}(f),\dots,L_{\text{train},m}(f)\right)$, and similarly with the validation, $L_{\text{val}}(f)$, and test data, $L_{\text{test}}(f)$. In this notation, the loss at the end of training can be expressed as $L_{(\cdot)}(f(\bm{p},T+1))$. When the $f$ being referred to is obvious, we simply write $L_{(\cdot)}^{t}(\bm{p})$, and for static mixtures we drop the superscript: $L_{(\cdot)}(\bm{p})$.
Data Mixing Problem. Given a set of data groups, an LM $f$ to train for $S$ steps with $N$ samples, and $T$ rounds of training (i.e., whether we use static or dynamic proportions), we aim to determine the $\bm{p}$ that minimizes the total test loss across groups: $\underset{\bm{p}\in\triangle^{T\times m}}{\text{minimize}}\sum_{i=1}^{m}L_{\text{test},i}^{T+1}(\bm{p}).$
This objective aims to produce a trained model that does well on many data groups, which can serve as a proxy for downstream performance. However, without assuming additional structure on $L^{T+1}(\bm{p})$, this problem can only be solved with a brute-force search over $\bm{p}$, which requires training many different models. In the next section, our LMO framework imposes a constraint on $L^{t+}(\bm{p})$ that allows many existing methods to be expressed as approaches to solving this problem.
@@ 3 A Unified Optimization Framework for Data Mixing
We introduce the LMO framework by stating the general optimization problem (Section 3.1). Then, we show how this framework can express several existing methods (Section 3.23.3), with a summary of our insights regarding these methods in Section 3.3.3.
3.1 Linear Mixing Optimization (LMO) Framework
The LMO framework consists of an optimization problem that is equivalent to the data mixing minimization problem (Section 2), subject to an additional constraint:
@@ 3.1 Linear Mixing Optimization (LMO) Framework
There are three components of this problem that need to be specified to yield a way to set $\bm{p}$: a) the parameterization of the mixing law ($T$, $\sigma$), b) the values of the parameters ($A^{t},b^{t},c^{t}$), and c) how to solve the problem. We express existing methods in LMO by specifying these components.
3.2 Preliminaries for unifying methods
We discuss preliminaries before presenting existing methods and explaining how they can be expressed in the LMO framework. First, we formally define what it means for a method to be expressed in the LMO framework. Then, we present a result that allows us to convert between linear dynamic mixing laws and a way to set $\bm{p}$, which we will to use to express online methods in our framework in Section 3.3.
@@ 3.2 Preliminaries for unifying methods
We say that a data mixing method can be expressed in the LMO framework if its exact algorithm—how it sets proportions $\bm{p}$ and trains model $f$ in terms of $\bm{p}$—can be equivalently constructed by specifying a mixing law and way of solving the LMO optimization problem.
This definition allows us to cast existing methods as a way of solving the LMO optimization problem based on how they set $\bm{p}$ and train according to $\bm{p}$, even if the methods themselves are not originally designed to minimize average test loss.
Converting mixing laws into update rules. When $T>1$, a natural way to solve the LMO optimization problem is via exponentiated gradient descent (EGD) [31, 5], which updates $p^{t}$ greedily while ensuring that it remains on the probability simplex. The following lemma presents the EGD update rule for the LMO optimization problem when $\sigma=\text{Id}$.
@@
This lemma shows how to adjust $p^{t}$ dynamically to solve the LMO optimization problem. Notably, this update rule is defined in terms of the mixing law parameters, $A^{t}$ and $b^{t}$. This gives us a way to convert between how a method sets $\bm{p}$ and the implicit assumption it makes in the mixing law.
@@
Table 1: Summary of how existing methods and Aioli are expressed in the LMO framework (1).
3.3 Unifying Existing Methods
We discuss four existing data mixing methods and express them as specific instances of the LMO framework. A summary of our insights is provided in Section 3.3.3 and Table 1. In Appendix B.1, we comment on how several other online and offline data mixing methods are related to our framework, and all proofs for this section are in Appendix B.2.
@@ 3.3.1 Offline methods
Data Mixing Laws (DML). Ye et al. [72] propose an offline method using a static mixing law ($T=1$): $L_{\text{val},i}(\bm{p})=c_{i}+b_{i}\exp(\sum_{j=1}^{m}-A_{ij}p_{j})$ for $i\in[m]$, with $A,b,c$ learned by sweeping training runs over static proportions ($\geq m+1$ runs to avoid being underdetermined). They select the proportion that minimizes the predicted validation loss. This law can be derived from (2) with $\sigma=\exp$, showing that LMO with a) log-linear static mixing law, b) fitted parameters, and c) direct computation of $\bm{p}$ can express DML.
@@ 3.3.2 Online Methods
We provide a colloquial description and an algorithmic description of the following three online methods. Then, in Theorem 1 we demonstrate how they all are expressed in LMO using a linear dynamic mixing law, the EGD update rule, and method-specific mixing law parameters.
Skill-It. Chen et al. [14] is an online method motivated by curriculum learning that dynamically adjusts mixture proportions. Data group interactions are expressed in a “skills graph,” where each edge denotes how much the loss on one group changes when trained on another. The skills graph is learned in advance using $m$ training runs and then used to update proportions $p^{t}$ throughout training.
@@
DoReMi. Xie et al. [70] is an online method that applies ideas from distributionally robust optimization to data mixing, where the training objective minimizes the worst-group excess loss over a model trained with stratified sampling. $p^{t}$ is updated dynamically to minimize this excess loss and then averaged for the final run. DoReMi requires two additional runs to learn a static $\bm{p}$.
@@
DoGE. Fan et al. [22] is an online method that solves a bi-level optimization problem in which $p^{t}$ is updated to minimize the average training loss at each step. By using a first-order Taylor approximation of the training loss, $p^{t}$ is updated using the gradient of each data group. The dynamic proportions are then averaged for the final run. DoGE requires one additional run to learn a static $\bm{p}$.
@@
Framework expression. All three online methods use an update rule $p_{j}^{t+1}\propto p_{j}^{t}\exp(\cdot)$, which is similar to (3). This provides intuition for our main theorem, which expresses these methods in LMO.
@@ 3.3.3 Summary of LMO Framework Insights
Table 1 summarizes how existing methods are expressed in the LMO framework. LMO reveals the assumptions each method makes through how the components of the framework are specified. First, all mixing laws are either linear or log-linear. Second, the mixing laws differ in the values of the parameters used. For example, Skill-It’s $A^{t}$ is the current loss times a static skills graph matrix, while DoReMi’s $A^{t}$ is diagonal. Third, offline mixing methods solve for $\bm{p}$ directly while online mixing methods use EGD, which uses a greedy approximation. If the mixing law and solving strategy assumptions hold true in practice, then the method yields optimal mixture proportions. In the next section, we study the fidelity of these assumptions.